OpenStudy (anonymous):

Solve the system of equations by graphing. Then classify the equations as dependent or independent and the system as consistent or inconsistent. 2x-y=8 8x-4y=8

OpenStudy (compassionate):

Oh - look who it is.

OpenStudy (anonymous):

yes and I'm still needing help. I understand slope, but I need a refresher over it. As for the rest, I don't know how to tell one from the other.

OpenStudy (compassionate):

So what exactly are you having trouble with? Is it solving the system, the depend-independent, or the consistant-inconsistant?

OpenStudy (compassionate):

[One Eternity Later...]

OpenStudy (anonymous):

Let's try this again.. it reloaded before I could post. I need help with all of it. I don't understand the last part of it at all. The consistent and dependent part. I need help remember the steps to solve the equation so I can graph it.

OpenStudy (compassionate):

2x-y=8 8x-4y=8 We want either the x or y term to cancel out. It's your choice. Look at both the equations. If we multiplied the first equation by -4, then we can cancel the Xs, right?

OpenStudy (e.mccormick):

"Solve the system of equations by graphing." If that is in written homework, Jewels89 will need the graph.

OpenStudy (compassionate):

In order to graph it you need to find your x and y points. ;)

OpenStudy (anonymous):

@e.mccormick its just a blank graph. I have to put the lines on it after I find the slope.

OpenStudy (anonymous):

and right. So multiply by -4?

OpenStudy (compassionate):

Correct. So do that

OpenStudy (e.mccormick):

Um.... They are two lines, no need to solve for the intersection. Just set \(x_1\) to 0, solve the \(y_1\) intercept. Then set \(y_1=0\) solve for \(x_1\). 2 points, draw line. Repeat for second line.

OpenStudy (anonymous):

\[-8x-4y=-12\]\[8x-4y=8\]\[-8y=-4\]\[y=0.5\]

OpenStudy (anonymous):

correct?

OpenStudy (e.mccormick):

Watch your sign on your y.

OpenStudy (e.mccormick):

\(2x-y=8\) \(-4\cdot [2x-y]=-4\cdot [8]\) Then what happens to the y?

OpenStudy (compassionate):

Huh?

OpenStudy (compassionate):

No, I continued to run into her threads. We were joking.